# Learning to Predict Structural Vibrations

### v3

@@ Abstract
In mechanical structures like airplanes, cars and houses, noise is generated and transmitted through vibrations. To take measures to reduce this noise, vibrations need to be simulated with expensive numerical computations. Surrogate deep learning models present a promising alternative to classical numerical simulations as they can be evaluated magnitudes faster, while trading-off accuracy. To quantify such trade-offs systematically and foster the development of methods, we present a benchmark on the task of predicting the vibration of harmonically excited plates. The benchmark features a total of 12000 plate geometries with varying forms of beadings, material and sizes with associated numerical solutions. To address the benchmark task, we propose a new network architecture, named Frequency-Query Operator, which is trained to map plate geometries to their vibration pattern given a specific excitation frequency. Applying principles from operator learning and implicit models for shape encoding, our approach effectively addresses the prediction of highly variable frequency response functions occurring in dynamic systems. To quantify the prediction quality, we introduce a set of evaluation metrics and evaluate the method on our vibrating-plates benchmark. Our method outperforms DeepONets, Fourier Neural Operators and more traditional neural network architectures. Code, dataset and visualizations: https://eckerlab.org/code/delden2023_plate
Keywords Machine Learning, Vibrations, Engineering, Acoustics, Vibroacoustics, Surrogate Modeling
@@ 1 Introduction
Figure 1: Left: We introduce a dataset of 12000 samples for predicting vibration patterns based on plate geometries. A harmonic force excites all plates, causing them to vibrate. The vibration patterns of the vibrating plates are obtained through numerical simulation. Diverse architectures are evaluated on the dataset. Right: Beading patterns are used in many vibrating technical systems. Here, on an oil filter, a washing machine and a disk drive. They increase the structural stiffness and alter the vibration.
@@
Vibrations of mechanical structures depend on the frequency of the excitation force (e.g. by the engine). A special case occurs when the excitation frequency matches an eigenfrequency of a given structure. In this case, the external force adds energy in phase with the structure’s natural vibration and amplifies the motion with each cycle. This continues until the energy added equals the energy lost due to damping, resulting in large vibration amplitudes. This effect is called resonance and leads to characteristic resonance peaks in the dynamic response of the system. At resonance frequencies, due to the higher vibration amplitudes, the noise emission is increased. A second distinctive feature of structural vibrations is the vibration pattern, i.e. the spatial field of vibration velocity amplitudes. With increasing frequency, these vibration patterns become more complex and exhibit more local maxima and minima (Figure 1, left) [2].
To reduce noise, the vibration patterns of a mechanical structure can be influenced through modifications to its design. One effective method is the placement of damping elements, that absorb vibrational energy and thereby reduce sound emission, but this adds weight and requires space. Another approach is introducing beadings, which are indentations in plate-like structures (Figure 1, right). Beadings increase the local stiffness of a structure, resulting in a shift in the structure’s eigenfrequencies and subsequent resonance peaks. When they are well-placed, beadings can reduce the vibration energy for a range of excitation frequencies by shifting the resonance peaks out of the range. Reducing the vibration energy for a specific range of frequencies is a goal in many applications, e.g. in automotive design, where a motor excites vibrations in a range of frequencies [3].
In this work, we focus on a crucial prerequisite for targeted modifications to a design: Computing its vibrational behavior. The finite element method (FEM) is an established approach for numerically solving partial differential equations. The geometry of a design is discretized into small elements and the solution of the PDE is approximated by simple functions, e.g. polynomial functions, defined on these elements. [4, 5]. This method enables the numerical simulation of vibration patterns, but is computationally expensive. With increasing frequency and decreasing wavelength, finer meshes are required to accurately resolve the vibrations. This leads to a high increase in computational load and limits the number of designs and value of the frequencies that can be evaluated. Deep learning surrogate models could potentially accelerate the evaluation of design candidates by several magnitudes.
@@
Our novel benchmark dataset consists of 12000 instances of an exemplary structural mechanical system, a plate excited by a harmonic force, and their numerically computed vibrations given a range of excitation frequencies. Given a plate instance, the task is to predict the vibration patterns and frequency response. We vary material properties of the plate as well as the geometry by adding beadings. Plates with beadings are abundant in technical systems (Figure 1, right). Plates are also often a component of more complex mechanical systems and their vibrational behavior on their own is similar to more complex systems [9, 10], making them a well-posed and scalable initial benchmark problem for deep learning methods.
To address the benchmark task, we propose a novel network architecture named Frequency-Query Operator (FQO). This model is trained to map plate geometries together with an excitation frequency query to the resulting vibration pattern. This approach is inspired by work on operator learning for predicting the solution to partial differential equations [11] and implicit models for shape representation [e.g. 12, 13, 14], which can both be evaluated at any point in the domain. In our case, this enables predictions for any excitation frequency, including those not seen during training. On our vibrating-plates benchmark, the proposed architecture is better able to predict the highly variable resonances occurring in vibration patterns and outperforms among others DeepONet [11] and Fourier Neural Operators [15].
@@ 2 Dataset and Benchmark Construction
2.1 Dataset
We introduce a dataset consisting of instances of aluminum plates and their vibration patterns. The plates are simply supported, i.e. the edges cannot move up and down, but they can rotate. The plate is excited by a harmonic point force at a fixed position near one corner of the plate with the excitation frequency varied between 1 and 300 Hz. While the specific setting in other mechanical engineering design tasks may differ, this setup functions as an exemplary engineering design problem. Analogous problems are the design of an air-conditioning enclosure [16], a washing machine [17] or parts of a car chassis [3]. Compared to these problems, our plate setup has two differences that allow for a comparatively easy experimental real world validation of the computed vibration patterns and do not change typical vibrational characteristics: First, exciting the plate with a point force near a corner avoiding symmetry lines is a typical choice for plate benchmarks, since it assures to excite all vibration modes of the plate. Second, the use of a simply supported plate mirrors e.g. a bonnet of a car that rests on the chassis. Other typical types of fixation include screws or welding.
In the following, we describe the specific quantity of interest of the vibration patterns, how the vibration patterns of the plate are obtained via numerical simulation and how the plate geometry and parameters are varied.
@@ 2.1 Dataset
Our benchmark is designed to cover a vibroacoustic engineering design problem. Therefore, the goal is to predict a quantity that best reflects the noise emitted by a mechanical structure. For a plate, a natural choice is the maximum velocity field $v_{z}(x,y|f)$ for a specific frequency $f$. Here, $v_{z}(x,y|f)$ represents the component of the velocity field orthogonal to the plate surface (in the following $\mathbf{V}(f)$ denotes the velocity field on the discrete grid). However, it includes too much spatial information to be easily interpretable. We consider the mean squared velocity as an integral quantity and represent it in a frequency response function $\mathcal{F}$, which is a function of the excitation frequency:
@@
The square velocity is proportional to the kinetic energy and is therefore closely related to how strongly the vibration couples into a surrounding fluid and can then be perceived as airborne sound. In the above expression, $A$ is the plate area for spatial averaging. The result is scaled by a reference value $r$ and converted to a decibel scale.
@@
We apply the finite element method to solve the shell formulation and simulate the vibrational behavior of the plate [4] (Figure 2). This involves partitioning the plate geometry into discrete elements and approximating the solution on these elements by simple ansatzfunctions. By choosing a sufficiently large number of elements, the solution converges to the exact solution of the model [23]. We discretize the plate with a regular grid of 81 x 121 nodes and use triangular elements in the domain to allow a flexible representation of beadings. The discretization is sufficient to resolve wave lengths in the plate structure, but limits the detail that can be represented with the beading patterns.
After discretizing the plate, the PDE is integrated over the elements and a linear system of equations is derived. This linear system describes the dynamics of the discretized structure and is solved with a direct solver. We perform the computations with a specialized FEM software for acoustics [24]. Further details on the setup and mechanical model are given in Appendix A.1.
Plate variations.
The plate instances are varied in two settings: For the V-5000 setting, we generate random beading patterns consisting of 1 - 3 lines and 0 - 2 ellipses. Also, the width of the beading-elements is randomly varied. The size of the plates as well as material parameters are fixed. For the G-5000 setting, we apply the same beading pattern variation and additionally vary the plate geometry (length, width and thickness) as well as one material parameter (damping loss factor). For each setting, 5000 instances for training and validation are generated. 1000 further instances are generated as a test set and are not used during training or to select a model. Further details are given in Appendix A.2.
@@
The mean plate design shows a close to uniform distribution, with an area left free from beadings at the plate’s edge (see Figure 3b). We find that the number of peaks corresponds with the beaded area: the greater the proportion of beaded area in a given plate, the fewer the peaks (see Figure 3a). This is due to additional beadings stiffening the plates, and it represents an interesting trait specific to our problem. The density of peaks is related to the frequency. As the frequency increases, so does the peak density. Starting from around 120 Hz the peak density plateaus (see Figure 3d). The average number of peaks in the G-5000 setting is smaller than in the V-5000 setting. This is influenced by the on average smaller plates being stiffer and therefore having less peaks in the frequency range (see Figure 3c).
@@ 2.2 Evaluation
We propose three complementary metrics to measure the quality of the frequency response predictions. To address numerical issues as well as facilitate an easier interpretation of the evaluation metrics, we normalize the frequency response and the velocity fields. To do this, we first take the log of the velocity fields, to align it with the dB-scale of the frequency response. Then, we subtract the mean per frequency over all samples (depicted in Figure 3b for frequency response) and then divide by the overall standard deviation across all frequencies and samples. Small changes in the beading pattern can cause frequency shifts, potentially pushing peaks out of the considered frequency band. To reduce the effect of such edge cases, we predict frequency responses between 1 and 300 Hz but evaluate on the frequency band between 1 and 250 Hz.
@@
To specifically address the prediction of resonance peaks, which are particularly relevant for noise emission, we introduce a third metric called peak frequency error. The metric answers two questions: (1) Does the predicted frequency response contain the same resonance peaks as the true response? (2) How far are corresponding ground truth and prediction peaks shifted against each other? To this end, we set up an algorithm that starts by detecting a set of peaks $K$ in the ground truth and a set of peaks $\hat{K}$ in the prediction using the find_peaks function in scipy [29] (examples in Appendix B). Then, we match these peaks pairwise using the Hungarian algorithm [30] based on the frequency distance. This allows us to determine the ratio between predicted and actual peaks $\frac{|\hat{K}|}{|K|}$ and $\frac{|K|}{|\hat{K}|}$. To equally penalize predicting too many and too few peaks we consider the minimum of both ratios: $\mathcal{E}_{\text{PEAKS}}=1-\min\{\frac{|\hat{K}|}{|K|},\frac{|K|}{|\hat{K}|}\}$.
@@ 3 Predicting Vibrations with Neural Networks
We propose a method to predict the frequency response, $\mathcal{F}_{\mathbf{g},\mathbf{m}}(f)$, for plates characterized by their geometry $\mathbf{g}$ (influenced by beading patterns) and material parameters $\mathbf{m}$ (height, width, thickness, damping loss factor). This process involves two steps: (1) First, the input $\mathbf{g}$ and $\mathbf{m}$ are encoded by an encoder $\Phi$. Because $\mathbf{g}$ is defined on a regular grid, standard image processing architectures are suitable. (2) Frequency response predictions are generated for specific excitation frequencies $f$ by a decoder $\Psi$ (Figure 4). The computation can then be expressed as:
$\Psi(\Phi(\mathbf{g},\mathbf{m}),f)=\mathcal{F}_{\mathbf{g},\mathbf{m}}(f)$ (2)
This problem formulation, training a neural network to predict a function and evaluating this function, given some input values, is a common paradigm in operator learning [11]. It allows for the evaluation of any frequency query $f$, even if it has not been part of the training data. In contrast, predicting frequencies on a fixed grid only allows the evaluation of those frequencies. This formulation shares similarities with implicit models, for instance by [13] in the context of 3d shape prediction. Based on this, we investigate the following central aspects of our architecture:
@@
Vibrations are dominated by resonance peaks, that are specific to a certain frequency. The exact frequency of resonances varies starkly between instances. An implicit or operator learning approach has been shown to be able to deal with high variation better in other contexts. In the context of vibration prediction, a frequency-query approach could be employed to generate predictions for one specific frequency.
@@
Image processing architectures based on convolutions encode local features. In contrast, vision transformers already have in early layers a global receptive field size. As vibrations are determined by the full geometry, we would expect vision transformers to perform better.
@@
We can train networks to either directly predict the aggregate frequency response $\mathcal{F}$ or to predict the velocity field $\mathbf{V}$ and compute $\mathcal{F}$ from $\mathbf{V}$ via Equation 1. For predicting the velocity field, much more training data is available, since it describes a field over the plate instead of the scalar frequency response. Most of this information is not represented in the frequency response.
In the following, we describe architectural variations explored for these aspects. See Appendix C for details on all architectures and specifics on the training procedure.

### v4

@@ Abstract
In mechanical structures like airplanes, cars and houses, noise is generated and transmitted through vibrations. To take measures to reduce this noise, vibrations need to be simulated with expensive numerical computations. Deep learning surrogate models present a promising alternative to classical numerical simulations as they can be evaluated magnitudes faster, while trading-off accuracy. To quantify such trade-offs systematically and foster the development of methods, we present a benchmark on the task of predicting the vibration of harmonically excited plates. The benchmark features a total of 12,000 plate geometries with varying forms of beadings, material, boundary conditions, load position and sizes with associated numerical solutions. To address the benchmark task, we propose a new network architecture, named Frequency-Query Operator, which predicts vibration patterns of plate geometries given a specific excitation frequency. Applying principles from operator learning and implicit models for shape encoding, our approach effectively addresses the prediction of highly variable frequency response functions occurring in dynamic systems. To quantify the prediction quality, we introduce a set of evaluation metrics and evaluate the method on our vibrating-plates benchmark. Our method outperforms DeepONets, Fourier Neural Operators and more traditional neural network architectures and can be used for design optimization. Code, dataset and visualizations: https://github.com/ecker-lab/Learning_Vibrating_Plates
@@ 1 Introduction
Figure 1: Left: We introduce the Vibrating Plates dataset of 12,000 samples for predicting vibration patterns based on plate geometries. A harmonic force excites the plates, causing them to vibrate. The vibration patterns of the plates are obtained through numerical simulation. Diverse architectures are evaluated on the dataset. Right: Beadings are indentations and used in many vibrating technical systems. Here, on an oil filter, a washing machine and a disk drive. They increase the structural stiffness and alter the vibration.
@@
Vibrations of mechanical structures depend on the frequency of the excitation force (e.g. by the engine). A special case occurs when the excitation frequency matches an eigenfrequency of a given structure. In this case, the external force adds energy in phase with the structure’s natural vibration and amplifies the motion with each cycle. This continues until the energy added equals the energy lost due to damping, resulting in large vibration amplitudes. This effect is called resonance and leads to characteristic resonance peaks in the dynamic response of the system. At resonance frequencies, due to the higher vibration amplitudes, more noise is emitted. A second distinctive feature of structural vibrations is the vibration pattern, i.e. the spatial field of vibration velocity amplitudes. With increasing frequency, these vibration patterns become more complex and exhibit more local maxima and minima (Figure 1, left) [2].
To reduce noise, the vibration patterns of a mechanical structure can be influenced through modifications to its design. One method is the placement of damping elements, that absorb vibrational energy and thereby reduce sound emission, but this adds weight and requires space. Another approach is introducing beadings, which are indentations in plate-like structures (Figure 1, right). Beadings increase the local stiffness of a structure, resulting in a shift in the structure’s eigenfrequencies and subsequent resonance peaks. When they are well-placed, beadings can reduce the vibration energy for a range of excitation frequencies by shifting the resonance peaks out of the range. Reducing the vibration energy for a specific range of frequencies is a goal in many applications, e.g. in automotive design, where a motor excites vibrations in a range of frequencies [3].
In this work, we focus on a crucial prerequisite for targeted modifications to a design: Computing its vibrational behavior. The finite element method (FEM) is an established approach for numerically solving partial differential equations. The geometry of a design is discretized into small elements and the solution of the PDE is approximated by simple functions, e.g. polynomial functions, defined on these elements. [4, 5]. This method enables the numerical simulation of vibration patterns, but is computationally expensive. With increasing frequency and decreasing wavelength, finer meshes are required to accurately resolve the vibrations. This leads to a high increase in computational load and limits the number of designs and value of the frequencies that can be evaluated. Deep learning surrogate models could accelerate the evaluation of design candidates by several magnitudes.
@@
Our novel benchmark dataset consists of 12,000 instances of an exemplary structural mechanical system, a plate excited by a harmonic force, and their numerically computed vibrations given a range of excitation frequencies. Given a plate instance, the task is to predict the vibration patterns and frequency response. We vary material properties and the boundary conditions of the plate as well as the geometry by adding beadings. Plates with beadings are abundant in technical systems (Figure 1, right). Plates are also often a component of more complex mechanical systems and their vibrational behavior on their own is similar to more complex systems [9, 10], making them a well-posed and scalable initial benchmark problem for deep learning methods.
To address the benchmark task, we propose a novel network architecture named Frequency-Query Operator (FQO). This model is trained to predict the resulting vibration pattern from plate geometries together with an excitation frequency query. This approach is inspired by work on operator learning for predicting the solution to partial differential equations [11] and implicit models for shape representation [e.g. 12, 13, 14], both techniques enable evaluating any point in the domain instead of a fixed grid. In our case, this enables predictions for any excitation frequency, including those not seen during training. On our vibrating-plates benchmark, the proposed FQO can accurately predict the highly variable resonances occurring in vibration patterns and outperforms DeepONet [11], Fourier Neural Operators [15] and other baselines.
@@ 2 Dataset and Benchmark Construction
2.1 Vibrating Plates Dataset
We introduce a dataset consisting of instances of aluminum plate geometries and their vibration patterns. The plates are simply supported, i.e. the edges cannot move up and down. Depending on the dataset setting, the rotational stiffness at the boundary is varied, which corresponds to free rotation or clamped edges. The plate is excited by a harmonic point force at varying positions with the excitation frequency varied between 1 and 300 Hz. While the specific setting in other mechanical engineering design tasks may differ, this setup functions as an exemplary engineering design problem. Analogous problems are the design of an air-conditioning enclosure [16], a washing machine [17] or parts of a car chassis [3]. Compared to these problems, our plate setup has two differences that allow for a comparatively easy experimental real world validation of the computed vibration patterns and do not change typical vibrational characteristics: First, exciting the plate with a point force is a common experimental setup, where a plate is excited via a shaker. Second, the condition of no rotational stiffness at the edges in comparison to clamped edges does not introduce additional uncertainty and parameters into the measurement and mirrors e.g. a bonnet of a car that rests on the chassis. Other typical types of fixation include screws or welding. In the following, we describe the specific quantity of interest of the vibration patterns, how the vibration patterns of the plate are obtained via numerical simulation and how the plate geometry and parameters are varied.
@@ 2.1 Vibrating Plates Dataset
Our benchmark is designed to address a vibroacoustic engineering design problem. Therefore, the goal is to predict a quantity that best reflects the noise emitted by a mechanical structure. For a plate, a natural choice is the maximum velocity field $v_{z}(x,y|f)$ for a specific frequency $f$. Here, $v_{z}(x,y|f)$ represents the component of the velocity field orthogonal to the plate surface (in the following $\mathbf{V}(f)$ denotes the velocity field on the discrete grid). This component closely relates to how much sound is radiated, but specific details about where the velocity on the plate is highest are superfluous. Therefore, we use the mean of the squared velocity as a more compact representation and express it in a frequency response function $\mathcal{F}$, which is a function of the excitation frequency:
@@
The square velocity is proportional to the kinetic energy and is therefore closely related to how strongly the vibration couples into a surrounding fluid and can then be perceived as airborne sound. In the above expression, $A$ is the plate area over which the velocity is averaged. The result is scaled by a reference value $r$ and converted to a decibel scale.
@@
We apply the finite element method to solve the shell formulation and simulate the vibrational behavior of the plate [4] (Figure 2). This involves partitioning the plate geometry into discrete elements and approximating the solution on these elements by simple ansatzfunctions. By choosing a sufficiently large number of elements, the solution converges to the exact solution of the model [23]. We discretize the plate with a regular grid and use triangular elements in the domain to allow a flexible representation of beadings. The discretization is sufficient to resolve wave lengths in the plate structure, but limits the detail that can be represented with the beading patterns. After discretizing the plate, the PDE is integrated over the elements and a linear system of equations is derived. This linear system describes the dynamics of the discretized structure and is solved with a direct solver. We perform the computations with a specialized FEM software for acoustics [24]. Further details on the setup and mechanical model are given in Appendix A.1.
Dataset variations.
The plate instances are varied in two settings: For the V-5000 setting, we generate random beading patterns consisting of 1 - 3 lines and 0 - 2 ellipses. Also, the width of the beading-elements is randomly varied. The size of the plates as well as material, boundary and loading parameters are fixed. For the G-5000 setting, we apply the same beading pattern variation and additionally vary the plate geometry (length, width and thickness) as well as the damping loss factor, rotational stiffness at the boundary and forcing position. For each setting, 5000 instances for training and validation are generated. 1000 further instances are generated as a test set and are not used during training or to select a model. Further details are given in Appendix A.2.
@@
The mean plate design shows a close to uniform distribution, with a margin at the plate’s edge (see Figure 3b). With a greater proportion of beaded area in a given plate, the number of peaks tends to decrease (see Figure 3a). This is due to additional beadings stiffening the plates, and it represents an interesting trait specific to our problem. The density of peaks is related to the frequency. As the frequency increases, so does the peak density. Starting from around 120 Hz the peak density plateaus (see Figure 3d). The average number of peaks in the G-5000 setting is smaller than in the V-5000 setting. This is influenced by the on average smaller plates being stiffer and therefore having less peaks in the frequency range (see Figure 3c).
@@ 2.2 Evaluation
Before computing our metrics, we perform the following preprocessing steps to address numerical issues as well as facilitate an easier interpretation of the evaluation metrics. We normalize the frequency response and the velocity fields. To do this, we first take the log of the velocity fields, to align it with the dB-scale of the frequency response. Then, we subtract the mean per frequency over all samples (depicted in Figure 3b for frequency response) and then divide by the overall standard deviation across all frequencies and samples. Small changes in the beading pattern can cause frequency shifts, potentially pushing peaks out of the considered frequency band. To reduce the effect of such edge cases, we predict frequency responses between 1 and 300 Hz but evaluate on the frequency band between 1 and 250 Hz.
We propose three complementary metrics to measure the quality of the frequency response predictions.
@@
To specifically address the prediction of resonance peaks, which are particularly relevant for noise emission, we introduce a third metric called peak frequency error. The metric answers two questions: (1) Does the predicted frequency response contain the same number of resonance peaks as the true response? (2) How far are corresponding ground truth and prediction peaks shifted against each other? To this end, we set up an algorithm that starts by detecting a set of peaks $K$ in the ground truth and a set of peaks $\hat{K}$ in the prediction using the find_peaks function in scipy [29] (examples in Appendix B). Then, we match these peaks pairwise using the Hungarian algorithm [30] based on the distance between the frequencies of the peaks $\mathcal{E}_{\text{F}}$. This allows us to determine the ratio between predicted and actual peaks $\frac{|\hat{K}|}{|K|}$ and $\frac{|K|}{|\hat{K}|}$. To equally penalize predicting too many and too few peaks we consider the minimum of both ratios: $\mathcal{E}_{\text{PEAKS}}=1-\min\{\frac{|\hat{K}|}{|K|},\frac{|K|}{|\hat{K}|}\}$.
@@ 3 Predicting Vibrations with Neural Networks
We propose a method to predict the frequency response, $\mathcal{F}_{\mathbf{g},\mathbf{m}}(f)$, for plates characterized by their geometry $\mathbf{g}$ (influenced by beading patterns) and scalar parameters $\mathbf{m}$ (height, width, thickness, damping loss factor, rotational stiffness at boundary, loading position). This process involves two steps: (1) First, the input $\mathbf{g}$ and $\mathbf{m}$ are encoded by an encoder $\Phi$. Because $\mathbf{g}$ is defined on a regular grid, standard image processing architectures are suitable. (2) Frequency response predictions are generated for specific excitation frequencies $f$ by a decoder $\Psi$ (Figure 4). The computation can then be expressed as:
$\Psi(\Phi(\mathbf{g},\mathbf{m}),f)=\hat{\mathcal{F}}_{\mathbf{g},\mathbf{m}}(f)$ (2)
This problem formulation, training a neural network to predict a function and evaluating this function, given some input values, is a common paradigm in operator learning [11]. It allows for the evaluation of any frequency query $f$, even if it has not been part of the training data. In contrast, predicting frequencies on a fixed grid only allows for the evaluation of those frequencies. This formulation shares similarities with implicit models, for instance by [13] in the context of 3d shape prediction. Based on this, we investigate the following central aspects of our architecture:
@@
Vibrations are dominated by resonance peaks at specific frequencies. The resonance frequencies vary strongly across instances. An implicit or operator learning approach has been shown to be able to deal with high variation better in other contexts. In the context of vibration prediction, a frequency-query approach could be employed to generate predictions for one specific frequency.
@@
Image processing architectures based on convolutions encode local features. In contrast, vision transformers have a global receptive field size from early layers. As vibrations are determined by the full geometry, we expect vision transformers to perform better.
@@
We can train networks to either directly predict the aggregate frequency response $\mathcal{F}$ or to predict the velocity field $\mathbf{V}$ and compute $\mathcal{F}$ from $\mathbf{V}$ via Equation 1. For predicting the velocity field, much richer training data is available, since it describes a field over the plate instead of the scalar frequency response. Most of this information is not represented in the frequency response.
In the following, we describe architectural variations explored for these aspects.